PLEASE HELP! SOLVE BY ADDITION METHOD. 3X-Y=11 AND 2X +5Y=13
Answer by bamigbe20(3)

Find x and y where 3X-Y=11 AND 2X +5Y=13
solution:
3X - Y = 11 ...................(1)
2X + 5Y = 13 ..................(2)
To find x, we must look for a way of eliminating y.
To do that we have to multiply equation 1 by 5, and equation 2 by 1.
i.e 5(3X - Y = 11)
1(2X + 5Y = 13)

it will then give us:
15X - 5Y = 55 ...............(3)
2X + 5Y = 13 ...............(4)
We add equation 3 and 4 together
..........................................
17X + 0 = 68
...........................................
so, 17X = 68; We then divide both sides by 17.
(17X)/17 = (68)/17
we now have X = 68/17
and X = 4.

To get Y, we substitute X = 4 into equation 1, considering equation 1 only.
3X - Y = 11
3(4) - Y = 11
12 - Y = 11
We collect like terms
-Y = 11 - 12
-Y = -1
then we divide both sides by -1

(-Y)/-1 = (-1)/-1

Y = 1
Therefore X = 4 and Y = 1